A) If a 2.5 L solution contains 0.141 moles of NaOH, what is the pH of the solution?

B) Calculate the percent ionization of a 0.44 M solution of acid with a pH of 3.81.
Answer #1

@ Molarity Of Naolt 0 141 mol =50648102m = 2.5 L 3 polt = -log [017] = -log (5064% 10% 1024 = 14-polt = 94-1024 = 12057 B THA